Graph the line that passes through the points (-3, -5) and (8, 6)?
What are the x- and the y-intercepts for this line?
x-intercept
a = ?
y-intercept
b = ?

Answer by mananth(16946)

plot the points (-3, -5) and (8, 6)
slope = 11/11=1
y=mx+b
6=1*8+b
b=-2
y=x-2

x intercept = 2
y intercept =-2